OpenStudy (anonymous):

Multiply: 2 the square root of 11 • 8 the square root of 22 a. 8 the square root of 11 b. 10 the square root of 33 c. 64 the square root of 22 d. 176 the square root of 2

OpenStudy (jack1):

\(\large 2\sqrt 11 \times 8\sqrt 22\)?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

I get to the point were it becomes 16 (sqrt) 242 then I dont know what to do

OpenStudy (jack1):

sqrt 11 * sqrt 22 = sqrt 11 * sqrt 11 * sqrt 2 = (sqrt 11)^2 * sqrt 2 = 11 * sqrt 2

OpenStudy (jack1):

so now add in the x 2 and x 8

OpenStudy (jack1):

11 * 2 * 8 * sqrt 2 = 176 * sqrt 2

OpenStudy (anonymous):

How did you get the sqrt 11 * sqrt 22

OpenStudy (jack1):

\(\Large 2\times \sqrt {11} \times 8 \times \sqrt {22}\) \(\Large 2\times 8 \times \sqrt {11} \times \sqrt {22}\) \(\Large 2\times 8 \times \sqrt {11} \times \sqrt {(2 \times 11)}\) \(\Large 2\times 8 \times \sqrt {11} \times \sqrt {2} \times \sqrt {11}\) \(\Large 2\times 8 \times \sqrt {11} \times \sqrt {11} \times \sqrt {2} \) \(\Large 2\times 8 \times 11 \times \sqrt {2} \)

OpenStudy (anonymous):

Oh so you were just simplifying the equation?

OpenStudy (jack1):

yeps ;)

OpenStudy (anonymous):

so that would be 176 (sqrt) 2

OpenStudy (jack1):

perfect

OpenStudy (anonymous):

Thank you so much! It make so much more sense now

OpenStudy (jack1):

all good, np ;)